• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name mcnulty

    Log entry time 10:47:01 on September27,2009

    Entry number 291913

    keyword=Target move error

    Target move error.  Clearing software record did not fix.  We rebooted the 
    control ioc's -- as instructed on the target troubleshooting guide.  During 
    this reboot we experienced a low temperature alarm (Target temp dropped to 
    17.2K).  The temp then stablized to 19K.  But target mover still not 
    working...we cleared the software record again.  Waited...and then clicked the 
    carbon hole position... and it worked.